# Omni-Router: Sharing Routing Decisions in Sparse Mixture-of-Experts for Speech Recognition

### v1

@@ VI-A Omni-router improves ASR
VI-B Omni-router learns specialized experts
To investigate whether the sharing of router decisions can lead to more specialized experts, we did a detailed analysis of the expert usage per token in the transformer-based ASR model with MoE layers, see Fig. 2. We visualize three randomly selected audio samples from the Librispeech dev-other dataset. Each example includes (from top to bottom): the mel-spectrogram, the expert assignments for a 4-expert Switch Transformer, and the expert assignments for a 4-expert Omni-router Transformer. Expert assignments are depicted across time steps on the horizontal axis and transformer layers on the vertical axis. Note that there is no specific meaning of the colors of the experts, and for Switch Transformer plots, colors are permuted to show the optimal mapping across layers.
We observe notable differences between the routing behaviors of the Switch Transformer and the Omni-router Transformer. The Omni-router Transformer demonstrates a markedly structured and coherent pattern of expert usage across both time and layers, whereas the Switch Transformer exhibits a comparatively scattered and less structured routing pattern. Specifically, the Omni-router’s routing decisions appear to form consistent temporal segments, suggesting meaningful coordination between layers. Moreover, there is a clear structure and usage of the same expert across the model’s depth, showing that the inductive bias of omni-router holds and provides consistent specialization. Furthermore, the Omni-router’s expert assignments clearly correlate with acoustic patterns in the mel-spectrogram. For instance, the green expert predominantly handles silent segments, whereas alternating stripes of orange, blue, and pink correspond to distinct speech regions. In contrast, the Switch Transformer displays more fragmented expert assignments, indicative of weaker inter-layer coordination and potentially less effective specialization.
@@ VI-B Omni-router learns specialized experts
To further quantify expert specialization in both models, we adopt the expert permutation method introduced by prior work [15]. Specifically, for each token, instead of using the top-1 expert selected by the router, we randomly select an expert with a certain probability $p$ and evaluate the impact on performance by monitoring changes in WER on the Librispeech dev-other set across varying permutation ratios. Our hypothesis is that the Omni-router Transformer ASR will exhibit a larger performance degradation (increased WER) than the Switch Transformer ASR, indicating greater expert specialization.
As illustrated in Fig. 3, we observe that as the permutation ratio $p$ increases, the Omni-router Transformer exhibits a higher sensitivity to incorrect expert selection, resulting in a more pronounced degradation in performance compared to the Switch Transformer ASR. This finding supports our hypothesis of enhanced expert specialization in the Omni-router model.
To understand how expert roles evolve between layers, we investigate the consistency of expert assignments in adjacent layers. Figure 4 visualizes this relationship for layers 14 and 15 using contingency tables, where expert indices have been reordered to show the optimal mapping. To quantify this effect across the entire network, Figure 5 plots the Cramér’s V correlation for all adjacent layer pairs. The Omni-router Transformer exhibits significantly stronger consistency in expert assignments, a trend that is particularly strong in deeper layers.
@@ VI-F Omni-router Transformer demonstrates better robustness
VII Conclusion

### v2

@@ V-D Evaluation
VI Results
VI-A Omni-router improves ASR
VI-B Why Does the Omni-router Transformer Outperform?
A central question raised by our experimental results is what factors contribute to the superior performance of the Omni-router Transformer. We hypothesize that the shared routing mechanism creates stable and structured ”pathways” for tokens through the network, enabling the model to develop more cooperative and specialized experts.
To investigate this hypothesis, we conducted a detailed analysis of expert utilization, as illustrated in Fig. 2. We visualize expert assignments across three randomly selected audio samples from the Librispeech dev-other dataset. Each sample visualization consists of (from top to bottom): the mel-spectrogram, the expert assignments for a 4-expert Switch Transformer, and the expert assignments for a 4-expert Omni-router Transformer. Expert assignments are plotted with time steps on the horizontal axis and transformer layers on the vertical axis. The colors representing experts have no inherent meaning, and for clarity in Switch Transformer plots, colors are permuted to optimally align across layers.
Our visualizations reveal significant differences in routing behavior between the two architectures. The Omni-router Transformer exhibits notably structured and coherent expert assignments across both temporal and depth dimensions, whereas the Switch Transformer demonstrates a more fragmented and less organized routing pattern. Specifically, the Omni-router’s routing creates consistent temporal segments, suggesting meaningful coordination and communication between layers. Furthermore, repeated utilization of the same experts across multiple layers indicates that the Omni-router’s inductive bias effectively promotes stable specialization. Crucially, expert assignments in the Omni-router Transformer clearly correlate with acoustic patterns in the mel-spectrogram. For example, the ”green” expert predominantly handles silent segments, while alternating patterns of orange, blue, and pink experts correspond clearly to distinct speech segments. In contrast, the Switch Transformer displays fragmented expert assignments, indicating weaker inter-layer coordination and suboptimal specialization.
To quantitatively substantiate these qualitative observations, we conducted analyses along three dimensions using the Librispeech dev-other dataset:
Enhanced Inter-Layer Cooperation: To understand how expert assignments evolve through layers, we examined the consistency of expert choices between adjacent layers. Fig.3 depicts this relationship using contingency tables for layers 14 and 15, with expert indices reordered for optimal alignment. Extending this analysis network-wide, Fig.4 shows Cramér’s V correlation values computed across all adjacent layer pairs. The results indicate that the Omni-router Transformer consistently achieves higher inter-layer expert assignment correlation, particularly pronounced in deeper network layers, reinforcing its stronger inter-layer cooperation.
Lower Expert Entropy: We further analyzed the confidence in routing decisions by evaluating expert entropy distributions across layers for the top 100 most frequently occurring tokens. The Omni-router Transformer consistently exhibits lower entropy compared to the Switch Transformer, indicating higher confidence and decisiveness in expert selection.
Figure 5: Expert entropy across layers of Switch Transformer and Omni-router Transformer.
Greater Expert Specialization: To rigorously quantify expert specialization, we adopted an expert permutation method proposed in prior research [15]. Specifically, we randomly permuted expert assignments for each token with varying probability $p$ and measured the resulting impact on ASR performance, quantified by WER degradation. We hypothesized that the Omni-router Transformer would exhibit a greater sensitivity to incorrect expert assignments, reflecting higher expert specialization.
@@ VI-B Why Does the Omni-router Transformer Outperform?
Fig. 6 supports our hypothesis, clearly illustrating that the Omni-router Transformer experiences a significantly larger performance degradation (increased WER) as the permutation ratio $p$ increases, compared to the Switch Transformer. This indicates a stronger reliance on accurate, specialized expert assignments and reinforces the conclusion that the Omni-router Transformer promotes enhanced expert specialization and cooperative behavior across its architecture.
@@ VI-F Omni-router Transformer demonstrates better robustness
VII Comparisons with Dense Models
To further showcase the effectiveness of Omni-router approach, we added a comparison of Omni-router ASR against dense baselines with similar total parameter counts. As shown in below table, the sparse Omni-router (246M total/140M active params) achieves a WER on par with both a dense model (210M) trained on the same data and the Whisper-small model (244M).
Dataset Omni-router (246M - 140M) Dense (210M) Whisper-small.en (244M)
AMI-IHM 17.8 17.8 17.6
Chime6 28.2 27.9 27.7
CommonVoice 16.4 17.6 15.2
Fleurs 8.4 8.8 7.6
Callhome 15.0 15.2 20
Switchboard 13.5 13.8 14.9
Librispeech (clean) 3.3 3.7 3.1
Librispeech (other) 7.3 8 7.4
WSJ (nov92) 3.7 4.1 3.5
Tedlium 4.2 4.3 4
Voxpopuli 8.5 9 8.2
VIII Conclusion
@@
Acknowledgment
The authors thank Dan Busbridge, Jason Ramapuram and Russ Webb for their insightful comments and Shuangfei Zhai for their assistance with the internal publication process.